OpenStudy (decentnabeel):

A truck tire has a diameter of 86cm and completes a revolution in 0.3 secod at top speed? how many meters does the tire travel in one minute when the truck is travelling at top speed

OpenStudy (decentnabeel):

@mathmale @mww @TheSmartOne

OpenStudy (decentnabeel):

@PizzaDelivery @Unavailabilityy

OpenStudy (decentnabeel):

@Hayhayz

OpenStudy (518nad):

every 3 secs is a revolution, diameter =86cm, circumferece = pi*diameter, circumference is the distance travelled every 3 secs the number of 3 second intervals in a min is 60/3 total distane travelled = 60/3 * circumference = 60/3 * 86cm*3.14

OpenStudy (decentnabeel):

not 3 second 0.3 second